Please solve the all the questions below. Thanks. Especially pay attention to 2nd question.

t, which type of proof is being used in each case to prove the theorem (A → C)? Last Line 겨 (p A -p) 겨 First Line a C b. C d. (some inference) C Construct a contrapositive proof of the following theorem. Indicate your assumptions and conclusion clearly 2. If you select three balls at random from a bag containing red balls and white balls, then at least two of the selected balls will have the same color 3. Write a direct proof of the following theorem. For all odd integers, n, (nA3 + 3) is even. Let P(x) be the proposition: Given positive integer x, for all positive real numbers a and b, (a + b)1x >= a^x + b^x 4. Prove P(0) using a vacuous proof. 5. Write a proof by contradiction of this theorem: Given positive integers x and y, if (x * y) = 1 then (x = y)
Answer #1

1.

(a) Proof by Contraposition

(b) Propositional Equivalence

(c) Vacuous proof

(d) Proof by rules of inference

2.

P is Select three balls at random from a bag containing red balls and white balls

Q is at least two of the selected balls will have same color

To prove: P -> Q

Contrapositive:

~Q -> ~P

~Q is having a single color.

Ball having single color implies that we can pick two red color balls or two white color balls. So, we are picking two balls from the bag.

~P is picking two balls from the bag which satisfies the condition.

Hence, the given statement is true by contraposition.

3.

Direct proof:

Given that n is odd

  • n=2*m+1

n3+3

= (2*m + 1)3 + 3

= 8*m3 + 1 + 6*m + 12*m2 + 3

= 8*m3 + 4 + 6*m + 12*m2

= 2* (4*m3 + 2 + 3*m + 6*m2)

Since, n3+3 is a multiple of 2, it must be even.

Hence proved.

4.

X is a positive integer.

Let’s say x is 1.

Left Hand Side = (a+b)x = (a+b)1 = a+b

Right Hand Side = ax + bx = a1 + b1= a+b

Here, Left Hand Side =Right Hand Side

Let’s say x is 2.

Left Hand Side = (a+b)x = (a+b)2 = a2+b2+2*a*b

Right Hand Side = ax + bx = a2 + b2

Since 2*a*b is an extra in left hand side, always it is greater.

Hence, Left Hand Side>=Right Hand Side

5.

X and y are positive integers.

To prove: if (x*y=1), then x and y are same and are equal to 1.

Proof by Contradiction:

Let’s say x or y are positive integers and one of them atleast is not equal to 1.

So, let’s say x is 1. However, y can’t be 1 as per the condition. So, let’s say y is 2.

Hence, x*y=2 which is not equal to 1.

So, if x*y=1, then x=y and x=y=1.

(Proved)